OpenStudy (anonymous):

Solve the Equation. Check the solution. -2/x+4=4/x+3

OpenStudy (mathmale):

Mind typing this equation in Equation Editor, for greater clarity? Otherwise the reader has to make assumptions about what you mean. For example, is that -2/x + 4, or is it -2/(x+4)? Once you've clarified this (preferably through using either parentheses or Equation Editor), others and I would be happy to help you solve this equation and to check your result.

OpenStudy (anonymous):

\[\frac{ -2 }{ x+4}=\frac{ 4 }{ x+3 }\]

OpenStudy (anonymous):

@mathmale There it is. Sorry for the confusion.

OpenStudy (anonymous):

Bring the RHS over to the LHS and then combine the LHS fractions.\[-\frac{2 (3 x+11)}{(x+3) (x+4)}=0 \] If the fraction equals zero then (3x+11)=0. Solve 3x+11=0 for x

OpenStudy (anonymous):

you'd end up with x=x with a remainder which doesn't fit any of my answers. @robtobey

OpenStudy (anonymous):

My answers are A. -13/6 B. -11 C.-8/3 D. -11/3

OpenStudy (anonymous):

3x+11=0 3x =-11 x =-11/3 D. is the answer.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

You are welcome and thank you for the medal.